OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (asnaseer):

lets call the smallest integer x. so what would the other two be in terms of x?

OpenStudy (raden):

x+1 and x+2

OpenStudy (asnaseer):

perfect, so now use the fact that the sum of these 3 numbers equals 267

OpenStudy (asnaseer):

oh - I thought @Guez replied here. @RadEn you should really let the asker try this.

OpenStudy (raden):

hehe, sorry...

OpenStudy (asnaseer):

np - it's just to guide the sker to the solution rather than just giving out an answer. that way you learn more. :)

OpenStudy (asnaseer):

*it's just good to...

OpenStudy (anonymous):

I dont understand what you mean when you ask what are the other two.

OpenStudy (asnaseer):

we set the smallest number equal x. now the question states that there are three consecutive integers. consecutive means "one after another" - e.g. 7, 8, 9 or 45, 46, 47 so if x is the smallest one, then the next one must be x+1. similarly, the third number must be (x+1)+1=x+2 hope that makes sense?

OpenStudy (anonymous):

So the largest intiger is (x + 267)+267

OpenStudy (anonymous):

*integer

OpenStudy (asnaseer):

no - what we did was to let the smallest one equal x, so we end up with these 3 numbers: smallest: x middle: x+1 largest: x+2 you are then told that their sum equals 267, so we can now say that: x + (x+1) + (x+2) = 267 solve this to find x. then x+2 will be your largest number.

OpenStudy (anonymous):

So the largest integer is 267?

OpenStudy (asnaseer):

look at what I typed. the largest number is (x+2). 267 is what you get when you add up all three numbers.

OpenStudy (anonymous):

So 2 is the largest number?

OpenStudy (asnaseer):

how did you get that>

OpenStudy (anonymous):

Im sorry your not making any sence, you wrote largest: x+2

OpenStudy (asnaseer):

we let x represent the smallest number. so x is an unknown quantity. do you understand this step?

OpenStudy (anonymous):

x = 90?

OpenStudy (asnaseer):

how did you arrive at that value?

OpenStudy (anonymous):

Expand the brackets X+X-1+X-2 =267 3X -3 = 267 3X = 270 X = 270/3 X = 90

OpenStudy (asnaseer):

almost - you made an error in the signs - you should be using X+1 and X+2 not X-1 and X-2

OpenStudy (asnaseer):

remember here we said x represents the smallest number

OpenStudy (anonymous):

So x =90

OpenStudy (raden):

i think we allow if done from the largest to the smallest

OpenStudy (asnaseer):

if you look back you will that I wrote the equation as: x + (x+1) + (x+2) = 267 this is what you need to solve

OpenStudy (asnaseer):

*will see